OpenStudy (caozeyuan):

Suppose a variable X is normally distributed with a standard deviation of 10. Given that 30.85% of the values of X are greater than 70, what is the mean of X?

OpenStudy (caozeyuan):

so Z=(70-mu)/10, and P(X>Z)=30.85%, solve for mu?

OpenStudy (knov):

any illustrations ?

OpenStudy (caozeyuan):

nope, thatsall of the question from our textbook

OpenStudy (knov):

k

OpenStudy (caozeyuan):

ok, so (70-mu)/10=invNorm(30.85%)

OpenStudy (caozeyuan):

sorry, invnorm(1-.3085)

OpenStudy (knov):

k, so here is a standard deviation formula